### v1

@@ Abstract
Engagement in virtual learning is crucial for a variety of factors including learner satisfaction, performance, and compliance with learning programs, but measuring it is a challenging task. There is therefore considerable interest in utilizing artificial intelligence and affective computing to measure engagement in natural settings as well as on a large scale. This paper introduces a novel, privacy-preserving method for engagement measurement from videos. It uses facial landmarks, which carry no personally identifiable information, extracted from videos via the MediaPipe deep learning solution. The extracted facial landmarks are fed to a Spatial-Temporal Graph Convolutional Network (ST-GCN) to output the engagement level of the learner in the video. To integrate the ordinal nature of the engagement variable into the training process, ST-GCNs undergo training in a novel ordinal learning framework based on transfer learning. Experimental results on two video student engagement measurement datasets show the superiority of the proposed method compared to previous methods with improved state-of-the-art on the EngageNet dataset with a %3.1 improvement in four-class engagement level classification accuracy and on the Online Student Engagement dataset with a %1.5 improvement in binary engagement classification accuracy. The relatively lightweight ST-GCN and its integration with the real-time MediaPipe deep learning solution make the proposed approach capable of being deployed on virtual learning platforms and measuring engagement in real time.
@@ 1 Introduction
Engagement is intricately linked to the interaction of students’ attention and interest within a specific context [10243609]. It emerges not just as a consequence of pre-existing interests but also has the capability to cultivate new interests through consistent attention [hidi2006four]. This aspect is extremely crucial in educational environments, where the development and maintenance of engagement play a pivotal role in facilitating content comprehension [10243609]. Nonetheless, the challenges of measuring and sustaining engagement demand significant mental and emotional exertion from instructors. The progression in remote sensing technologies, Artificial Intelligence (AI), and affective computing holds promise for better measurement of engagement in diverse learning settings, including remote and virtual learning programs. Utilizing technologies such as facial expression recognition [liu2022graph], eye gaze tracking [wood2015rendering], and physiological stress detection [gedam2021review], instructors can more effectively monitor and enhance student engagement. This paper covers the development of AI algorithms for automated measurement of student engagement in virtual learning programs, taking into account the definition of engagement in educational psychology and its measurement methodologies.
Fredricks et al. [fredricks2004school] introduced three components of engagement: affective, behavioral, and cognitive engagement. Affective engagement relates to the emotions and attitudes a student has towards a task or its context, for instance, interest in a subject or feelings of excitement or confusion about a topic during a learning class [d2012dynamics]. Behavioral engagement involves the student’s active participation in a task, with indicators including looking at a computer screen (an on-task behavior) or playing with a mobile phone (an off-task behavior) during a learning class [ocumpaugh2015baker]. Cognitive engagement pertains to the student’s commitment to learning and their readiness to embrace challenges, encompassing cognitive outcomes such as recall, learning, and deep understanding [10243609].
Booth et al. [10243609] identified that the primary signals used to measure engagement components include facial, eye gaze, and body pose features, heart rate, Electro-Encephalo-Graphy, audio, and computer logs. Given that many of these signals can be captured by cameras and considering the widespread presence of cameras in personal computers and mobile phones—the devices commonly used for participation in virtual learning programs—video data has emerged as the predominant data modality for automatic AI-driven engagement measurement [karimah2022automatic, dewan2019engagement]. Consequently, most AI algorithms for engagement measurement have been designed and developed to analyze video data and infer the level of student engagement from the video [karimah2022automatic, dewan2019engagement, khan2022inconsistencies].
Engagement measurement methods are typically divided into end-to-end or feature-based categories [karimah2022automatic, dewan2019engagement, khan2022inconsistencies]. While feature-based approaches usually surpass end-to-end methods (that process raw videos) in performance [abedi2023affect, karimah2022automatic], they are constrained by the need to extract and identify the most effective features, often through trial and error. Typically, features from various modalities must be extracted from videos using multiple computer vision algorithms or deep neural networks. The extracted features are then analyzed by other deep neural networks or traditional machine learning methods to determine engagement [abedi2023affect, karimah2022automatic]. This requirement renders multi-modal feature-based approaches computationally heavy and challenging for practical implementation in real-world virtual learning platforms. These computationally intensive approaches render them impractical for local devices and necessitate the transfer of privacy-sensitive raw video from local devices to cloud servers to perform automated engagement measurement. Compared to raw videos, low-dimensional landmark information extracted from videos, such as facial landmarks, hand landmarks, and body joints [lugaresi2019mediapipe], offers not only more compact data but also encapsulates essential discriminative geometric features necessary for human affect [wei2023geometric] and behavior analysis [10144079], including engagement indicators. Most importantly, this geometric information is devoid of personally identifiable information [mishra2022skeletal].
@@
  • This is the first instance in the field of video-based engagement measurement [karimah2022automatic, dewan2019engagement, khan2022inconsistencies, chen2019faceengage, wu2020advanced, ma2021automatic, copur2022engagement, abedi2023affect, abedi2023detecting, thomas2018predicting, thomas2022automatic, gupta2016daisee, zhang2019novel, abedi2021improving, mehta2022three, fwa2022fine] where facial landmarks extracted from videos, as the unique data modality, are analyzed using Spatial-Temporal Graph Convolutional Networks (ST-GCNs) [yan2018spatial] to infer the engagement level in the video;
@@
  • Extensive experiments conducted on two video-based engagement measurement datasets demonstrate the superiority of the proposed method over previous methods, achieving an improved state-of-the-art in engagement level classification accuracy.
The structure of this paper is organized as follows. Section 2 offers an overview of the relevant literature. This is succeeded by Section 3, which details the proposed methodology. Following this, Section 4 describes the experimental setup and discusses the results obtained with the proposed method. Lastly, Section 5 concludes the paper and proposes directions for future research.
@@ 2 Related Work
In line with the contributions of the proposed work in this paper, the literature review presented in this section serves a dual purpose: it discusses previous video-based engagement measurement techniques and explores the methods previously utilized variants of graph convolutional networks for the purpose of facial expression and affect analysis.
@@ 2.1 Engagement Measurement
Karimah et al. [karimah2022automatic] conducted a systematic review and explored the measurement of student engagement in virtual learning environments. They found that most engagement measurement methods primarily focused on the affective and behavioral components of engagement. These methods commonly employed datasets annotated by external observers [khan2022inconsistencies] to train feature-based and end-to-end machine-learning and deep-learning models. In the following, some of the relevant feature-based and end-to-end works on video-based engagement measurement are discussed.
@@
In the domain of end-to-end engagement measurement techniques, deep neural networks analyze consecutive raw video frames to output the engagement level of the student in the video. These methods do not employ the extraction of handcrafted features from the videos; instead, the network is adept at autonomously learning to extract the most salient features directly from the videos, utilizing consecutive convolutional layers. The deep neural networks implemented in these end-to-end approaches include networks capable of video analysis, such as 3D Convolutional Neural Networks (3D CNNs), as well as combinations of 2D CNNs with sequential neural networks. For engagement measurement with 3D video input, networks such as C3D [gupta2016daisee, abedi2021improving], Inflated 3D CNN (I3D) [zhang2019novel], 3D DenseNet [mehta2022three], and Video Transformer [ai2022class] have been utilized. In subsequent methods, 2D CNNs and sequential models are jointly trained to automatically extract visual features from raw video frames, analyze the sequences of these features, and output the engagement level. Examples of this approach include the integration of 2D CNN with Long Short-Term Memory (LSTM) [gupta2016daisee], a combination of 2D Residual Neural Network (ResNet) and Temporal Convolutional Network (TCN) [abedi2021improving], ResNet paired with LSTM equipped with global attention [liao2021deep], and EfficientNet combined with LSTM and TCN [selim2022students].
@@
The process of measuring engagement through feature-based techniques is divided into two stages. The first stage involves extracting behavioral and affective features from video frames, which depends on either specific domain knowledge or the use of pre-trained models for facial embedding extraction. Recent examples of pre-trained models used for extracting facial embedding features from videos for engagement measurement include Masked Autoencoder for facial video Representation LearnINg (MARLIN) [cai2023marlin] utilized by Singh et al. [engagenet] and Emotion Face Alignment Network (EmoFAN) [toisoul2021estimation] utilized by Abedi et al. [abedi2023affect]. OpenFace [baltrusaitis2018openface] is notably recognized for its effectiveness in extracting features associated with facial Action Units (AUs), eye movements, gaze direction, and the positioning of the head. OpenFace [baltrusaitis2018openface] has been widely used in feature-based engagement measurement techniques. The extracted features are then combined in various ways to create sequences of feature vectors for the video. These feature vector sequences, taken from consecutive video frames, form multivariate time series data. To analyze the multivariate time series, machine learning and deep learning models such as Bag-of-Words (BoW) models [abedi2023bag], variations of Recurrent Neural Networks (RNNs) [abedi2023detecting], TCNs [abedi2023affect], Transformers [engagenet], or ensemble models [tian2023predicting] are employed. These models are trained on the sequences of features to derive levels of engagement [copur2022engagement, whitehill2014faces, liao2021deep, niu2018automatic, thomas2018predicting, huang2019fine, fedotov2018multimodal, chen2019faceengage, booth2017toward, kaur2018prediction, wu2020advanced, ma2021automatic].
Prior studies have demonstrated the advantages of feature-based methods over end-to-end approaches in measuring engagement. Yet, feature-based methods are constrained by the need to extract and pinpoint the most discriminative features, often through a process of trial and error. Additionally, they necessitate the segmentation of videos into multiple segments (comprising multiple frames) and the extraction of diverse multi-modal features from all video segments using multiple deep neural networks. For example, techniques employing OpenFace 2.0 [baltrusaitis2018openface] for extracting features such as eye gaze, head pose, and Action Units (AUs) involve the use of multiple complex neural networks [wood2015rendering, zadeh2017convolutional, zhang2016joint]. This makes the methods time-intensive during the inference phase, as the features extracted must be further analyzed by another deep neural network to make a final inference regarding engagement [engagenet].
@@ 2.2 Graph-based Facial Affect and Expression Analysis
Liu et al. [liu2022graph] conducted a comprehensive review of the literature on graph-based methods for facial affect analysis. These methods typically take an image or a sequence of images as input and produce an affect classification or regression as output. Based on their review, Liu et al. [liu2022graph] proposed a pipeline for graph-based facial affect analysis, which includes (i) face preprocessing, (ii) graph-based affective representation, and (iii) graph-based relational reasoning. The preprocessing involves necessary steps such as face detection and registration. Graph-based affective representation involves defining the structure of the graph, i.e., nodes and edges. The graph structure can be spatial, or spatiotemporal when the input data is still images, or video frames, respectively. The graph structure can be at landmark-level, region-level, or AU-level with nodes representing facial landmarks, facial regions of interest, or facial AUs, respectively. In the relational reasoning step, the edges, nodes, their interrelations, and temporal dynamics are analyzed through relational reasoning machine-/deep-learning models to make an inference regarding affect. The machine-/deep-learning models used to analyze graph data include dynamic Bayesian networks, RNNs, CNNs, fully-connected neural networks, and non-temporal and temporal graph neural networks [liu2022graph].
@@
The facial expression recognition method proposed by Zhou et al. [9191181] is among the methods where affective representation was performed by a spatiotemporal landmark-level and region-level graph. The method utilized thirty-four facial landmarks located around the eyes, lips, and cheeks, and the connections among these landmarks formed the intra-frame graph. These intra-frame connections were defined manually. To establish the inter-frame connections, each node in a frame was linked to its corresponding node in the subsequent frames. Two parallel ST-GCNs with analogous structures were trained: one on the x and y coordinates of the nodes, and the other on the histogram of orientation features derived from regions surrounding each node. The outputs from these two ST-GCNs were concatenated and then processed through a fully connected network for recognizing facial expressions in the input video. Among the disadvantages of this method are the independent training of the above-mentioned two ST-GCNs rather than joint learning and the manual definition of nodes and edges rather than allowing the ST-GCNs to learn the most relevant nodes and edges. Nodes were selected from Dlib’s facial landmarks [king_dlib] and additional nodes were created on the cheeks and between the eyes according to Dlib’s facial landmarks. The selection and creation of nodes were based on their relevance to certain AUs associated with the facial expressions that should be recognized.
@@
Wei et al. [wei2023geometric] introduced a graph-based method for micro-expression recognition in video. The introduced approach included a dual-stream ST-GCN, focusing on the $x$ and $y$ coordinates of facial landmarks and the distance and angles between adjacent facial landmarks. Recognizing the association between AUs and micro-expressions, they integrated an AU-specific loss function into their neural network’s training process. Their methodology employed three different sets of facial landmarks as graph nodes, comprising sets with 14, 31, and Dlib’s [king_dlib] 68 facial landmarks. Notably, the set with 68 landmarks was the only one to include landmarks along the jawline. The experiments demonstrated the best results of micro-expression recognition when 14 facial landmark sets were used.
@@
Leong et al. [fwa2022fine] introduced a method for detecting academic emotions through the use of facial landmarks and the roll, pitch, and yaw of head pose, utilizing spatial-temporal graph attention networks thereafter. Facial regions are cropped from video frames, and facial landmarks from these images are extracted, retaining those around the brow, eye, nose, and mouth, while excluding those that outline the face’s outer shape, with the reasoning being they lack correlation with affective states. These features are then inputted into spatial-temporal graph attention networks to determine academic emotions. A notable limitation of this approach is its reliance on multiple deep neural networks for extracting features, i.e., facial landmarks and head pose. The method achieved lower academic emotion detection accuracy when compared to prior feature-based methods.
@@ 2.3 Discussion
As reviewed in this section, while some methods have utilized facial landmarks and ST-GCNs for video-based facial affect and expression recognition, there has been no exploration in using facial landmarks or ST-GCNs specifically for video-based engagement measurement. It is important to note the fundamental differences between engagement and facial affect and expression making their measurement different. First, engagement is a multi-component state comprising behavioral and cognitive components in addition to its affective component. To illustrate, key indicators of behavioral engagement, such as being off-task or on-task, are determined by head pose, eye gaze, and blink rate. These indicators, and therefore engagement, cannot be effectively measured using methods designed solely for facial affect analysis. Second, engagement is not a constant state; it varies over time and should be measured at specific time resolutions where it remains stable and quantifiable. The ideal duration for measuring engagement is between 10 and 40 seconds, which is longer than the time resolution for facial affect analysis which is sometimes at frame level. Third, engagement measurement could involve recognizing an ordinal variable that indicates various levels of engagement, which differs from facial expression recognition, where the goal is to identify a categorical variable with no inherent order.
The limitations of existing engagement measurement approaches, discussed at the end of subsection 2.1, along with the key differences outlined above between facial affect analysis and engagement measurement, highlight a gap in the field of engagement measurement. To address this gap, we propose a simple yet effective engagement measurement technique. This method is also privacy-preserving and demonstrates improvements over existing methods on two video-based engagement measurement datasets.

### v2

@@ Abstract
Engagement in virtual learning is crucial for a variety of factors including student satisfaction, performance, and compliance with learning programs, but measuring it is a challenging task. There is therefore considerable interest in utilizing artificial intelligence and affective computing to measure engagement in natural settings as well as on a large scale. This paper introduces a novel, privacy-preserving method for engagement measurement from videos. It uses facial landmarks, which carry no personally identifiable information, extracted from videos via the MediaPipe deep learning solution. The extracted facial landmarks are fed to Spatial-Temporal Graph Convolutional Networks (ST-GCNs) to output the engagement level of the student in the video. To integrate the ordinal nature of the engagement variable into the training process, ST-GCNs undergo training in a novel ordinal learning framework based on transfer learning. Experimental results on two video student engagement measurement datasets show the superiority of the proposed method compared to previous methods with improved state-of-the-art on the EngageNet dataset with a 3.1% improvement in four-class engagement level classification accuracy and on the Online Student Engagement dataset with a 1.5% improvement in binary engagement classification accuracy. Gradient-weighted Class Activation Mapping (Grad-CAM) was applied to the developed ST-GCNs to interpret the engagement measurements obtained by the proposed method in both the spatial and temporal domains. The relatively lightweight and fast ST-GCN and its integration with the real-time MediaPipe make the proposed approach capable of being deployed on virtual learning platforms and measuring engagement in real-time.
@@ 1 Introduction
Engagement is key in education, blending students’ attention and interest within a learning context [1]. It not only stems from existing interests but also fosters new ones through sustained attention [2], essential for content comprehension and engagement development [1]. However, measuring and upholding engagement is challenging, and demands significant effort from educators. The advent of remote sensing, Artificial Intelligence (AI), and affective computing offers new avenues for accurately measuring engagement across various learning environments, including virtual learning platforms. Technologies such as facial expression recognition [3] and eye gaze tracking [4] enable more precise monitoring and enhancement of student engagement. This paper explores the development of AI algorithms for automated engagement measurement in virtual learning, taking into account the definition of engagement in educational psychology and its measurement methodologies.
Fredricks et al. [5] defined engagement through affective, behavioral, and cognitive components. Affective engagement relates to students’ emotions and attitudes towards their tasks, including interest or feelings during a learning class [6]. Behavioral engagement refers to active participation, such as focusing on a computer screen or not being distracted by a phone [7]. Cognitive engagement deals with a student’s dedication to learning and tackling challenges, affecting outcomes such as recall and understanding [1]. Booth et al. [1] identified that engagement is measured through various signals such as facial expressions, eye movements, posture, heart rate, brain activity, audio, and digital interactions. Cameras, prevalent in devices for online learning, make video a key data modality for AI-based engagement measurement [8, 9]. Therefore, AI techniques predominantly focus on video to measure student engagement [8, 9, 10].
AI-driven engagement measurement methods are divided into end-to-end and feature-based approaches. While feature-based methods generally outperform end-to-end approaches, which process raw videos, they require extensive identification of effective features through trial and error [8, 9, 10]. Extracting multi-modal features from videos with multiple computer vision algorithms or neural networks [11, 12], and their subsequent analysis by deep neural networks to measure engagement, render these methods computationally demanding [13, 8]. Such computational requirements limit their use on local devices and necessitate the transfer of privacy-sensitive video data to cloud servers for engagement measurement. In contrast, extracting low-dimensional landmark information, such as facial and hand landmarks, not only provides more compact data but also captures essential geometric features for affect and behavior analysis, including engagement, without personal identifiers [14, 15, 16]. Previous feature-based engagement measurement approaches extracted features such as head pose, facial Action Units (AUs), iris and gaze features, and affect and emotion features, which are indicators of the behavioral and affective components of engagement [8, 9, 10]. The literature has demonstrated the success of facial landmarks in capturing these aforementioned features [17, 18, 19, 20, 3]. A model such as Spatial-Temporal Graph Convolutional Networks (ST-GCNs) [21], capable of analyzing spatial-temporal facial landmarks and learning these aforementioned features inherently, can reduce the need for raw facial videos. An approach based on facial landmarks prioritizes privacy and reduces computational demands, making it more practical for real-time engagement measurement.
@@
  • This marks the first instance in video-based engagement measurement [8, 9, 22, 23, 13, 24, 25] where facial landmarks, as the single data modality extracted from videos, are analyzed through ST-GCNs [21] to infer the engagement level in the video;
@@
  • Extensive experiments conducted on two video-based engagement measurement datasets demonstrate the superiority of the proposed method over previous methods, achieving an improved state-of-the-art in engagement level classification accuracy. For explainability, Gradient-weighted Class Activation Mapping (Grad-CAM) is applied to the developed ST-GCNs to interpret the engagement measurements obtained by the proposed method in both spatial and temporal domains.
@@ 2 Related Work
The literature review in this paper serves two main purposes: discussing past video-based engagement measurement techniques and examining the use of graph convolutional networks for facial expression and affect analysis.
@@ 2.1 Engagement Measurement
Karimah et al. [8] conducted a systematic review on measuring student engagement in virtual learning environments, revealing a focus on affective and behavioral components of engagement. Most methods utilize datasets annotated by external observers [10] to train both feature-based and end-to-end models. In the following, some of the relevant feature-based and end-to-end works on video-based engagement measurement are discussed.
@@
In the domain of end-to-end engagement measurement techniques, deep neural networks analyze consecutive raw video frames to output the engagement level of the student in the video. These methods do not employ the extraction of handcrafted features from the videos; instead, the network is adept at autonomously learning to extract the most useful features directly from the videos, utilizing consecutive convolutional layers. The deep neural networks implemented in these end-to-end approaches include networks capable of video analysis, such as 3D Convolutional Neural Networks (3D CNNs) and Video Transformers [26, 27, 28], as well as combinations of 2D CNNs with sequential neural networks such as Long Short-Term Memory (LSTM) and Temporal Convolutional Network (TCN) [26, 27, 29, 30].
@@
The process of measuring engagement through feature-based techniques involves two stages. Initially, behavioral and affective features are extracted from video frames, relying on either domain-specific knowledge or pre-trained models for facial embedding extraction. OpenFace [31] is notably effective for its comprehensive feature extraction capabilities, including AUs, eye movements, gaze direction, and head positioning, and is widely applied in engagement measurement [13, 32, 33, 12]. Examples of facial embedding include the Masked Autoencoder for facial video Representation LearnINg (MARLIN) [11], utilized by Singh et al. [12], and the Emotion Face Alignment Network (EmoFAN) [19], used by Abedi et al. [13]. Subsequently, to analyze these extracted features and infer engagement, various machine learning and deep learning models are employed, such as Bag-of-Words (BoW) [32], Recurrent Neural Network (RNN) variations [24], Temporal Convolutional Networks (TCNs) [13, 33], Transformers [12, 34], and ensemble models [35].
@@ 2.2 Graph-based Facial Affect and Expression Analysis
Liu et al. [3] conducted a comprehensive review of the literature on graph-based methods for facial affect analysis. These methods typically take an image or a sequence of images as input and produce an affect classification or regression as output. Based on their review, Liu et al. [3] proposed a pipeline for graph-based facial affect analysis, which includes (i) face preprocessing, (ii) graph-based affective representation, and (iii) graph-based relational reasoning. Preprocessing involves steps such as face detection and registration. Graph-based affective representation involves defining the structure of the graph, i.e., nodes and edges. The graph structure can be spatial or spatiotemporal depending on whether the input data is still images or videos. The graph structure can be at the landmark level, region level, or AU level, with nodes representing facial landmarks, facial regions of interest, or facial AUs, respectively. In the relational reasoning step, the edges, nodes, their interrelations, and temporal dynamics are analyzed through relational reasoning machine-learning or deep-learning models to make inferences regarding affect. Models used to analyze graph data include dynamic Bayesian networks, RNNs, CNNs, fully-connected neural networks, and non-temporal and temporal graph neural networks [3].
@@
Zhou et al. [36] proposed a facial expression recognition method based on spatiotemporal landmark-level and region-level graphs. The intra-frame graph is formed by the connections among thirty-four facial landmarks located around the eyes, lips, and cheeks. The definition of these intra-frame connections was done manually. Inter-frame connections were established by linking each node in one frame to its corresponding node in the following frame. Two parallel ST-GCNs with analogous structures were trained; one on the nodes’ $x$- and $y$-coordinates, and another on their histogram of orientation features. ST-GCNs’ outputs were concatenated and processed by a fully connected network to identify facial expressions. This method’s disadvantages include independent training of the two ST-GCNs rather than joint learning, and manual definition of nodes and edges.
@@
Wei et al. [15] presented a graph-based method for micro-expression recognition in video. The method included a dual-stream ST-GCN, focusing on the $x$ and $y$ coordinates of facial landmarks and the distance and angles between adjacent facial landmarks. An AU-specific loss function was incorporated into the neural network’s training process in order to incorporate the association between AUs and micro-expressions. Their methodology employed three different sets of facial landmarks as graph nodes, comprising sets with 14, 31, and Dlib’s [37] 68 facial landmarks. Notably, the set with 68 landmarks was the only one to include landmarks along the jawline. The experiments demonstrated the best results for micro-expression recognition when 14 facial landmark sets were used.
@@
Leong et al. [25] introduced a method employing spatial-temporal graph attention networks to analyze facial landmarks and head poses, aimed at identifying academic emotions. The facial landmarks used were those around the eyebrow, eye, nose, and mouth excluding those that outline the face’s outer shape, with the reasoning being they lack correlation with affective states. A notable limitation of this approach is its reliance on multiple deep neural networks for extracting features, i.e., facial landmarks and head pose. The method achieved lower academic emotion detection accuracy when compared to prior feature-based methods [13].
@@ 2.3 Discussion
As reviewed in this section, while some methods have used facial landmarks and ST-GCNs to recognize facial affect and expression in videos, there has been no exploration of the use of these methods to measure engagement. The fundamental differences between engagement and facial affect and expression make their measurement different. First, engagement is a multi-component state comprising behavioral, affective, and cognitive components. To illustrate, key indicators of behavioral engagement, such as being off-task or on-task, are determined by head pose, eye gaze, and blink rate. These indicators, and therefore engagement, cannot be effectively measured using methods designed solely for facial affect analysis. Second, engagement is not a constant state; it varies over time and should be measured at specific time resolutions where it remains stable and quantifiable. An ideal measurement duration for engagement is between ten and forty seconds, which is longer than the time resolution for facial affect analysis, which sometimes occurs at the frame level. Third, engagement measurement could involve recognizing an ordinal variable that indicates levels of engagement, as opposed to facial expression recognition, which identifies categorical variables without inherent order.
Existing engagement measurement approaches face limitations due to the necessity of employing multiple deep neural networks for the extraction and analysis of multi-modal features. Coupled with the significant differences between facial affect analysis and engagement measurement as outlined above, these limitations underscore a gap in the field. In response, we introduce a straightforward yet effective method for engagement measurement. This method is lightweight and fast, preserving privacy while also demonstrating improvements over current methodologies across two video-based engagement measurement datasets.